OpenStudy (anonymous):

The equation of a hyperbola is shown. x^2 / 25 - y^2 / 4 = 1 can someone help me and i need to show my work What are the equations of the asymptotes? Option A: y=2/5 x and y=-2/5 x Option B: y= 4/5 x and y=-4/5 x Option C: y=5/4 x and y=-5/4 x Option D: y=5/2 x and y=-5/2 x

OpenStudy (anonymous):

Option A : y=2/5x and y = -2/5x

OpenStudy (anonymous):

I agree @shivshilpa